OpenStudy (anonymous):

what is the equation of the line that passes through the points (3,2)and has a slope of 4

OpenStudy (anonymous):

are you familiar with y=mx+c

OpenStudy (anonymous):

Yes but I get confused

OpenStudy (anonymous):

alright. do you know what m represents?

OpenStudy (anonymous):

Slope

OpenStudy (anonymous):

aren't we given the slope though?

OpenStudy (anonymous):

so automatically we can say y=4x+c

OpenStudy (anonymous):

follow?

OpenStudy (anonymous):

Yes but then C is what?

OpenStudy (anonymous):

well 'c' is the y-intercept of the graph right?

OpenStudy (anonymous):

what other information have you been given?

OpenStudy (anonymous):

@chris00 cant she use \[y - y_{1}=m(x-x_{1})\] and just solve for y?

OpenStudy (anonymous):

you need two points though...your only given one. and that only gives the slope which we are have already been given in the question

OpenStudy (anonymous):

The other information are the answers

OpenStudy (anonymous):

\[y - 2 = 4(x-3)\] solve for y you get a \[y = mx + b\] and you will have your equation

OpenStudy (anonymous):

i never learnt it that way. when you know y=4x+c substitute (3,2) into x any y 2=4(3)+c c=-10 therefore your equation of your line is y=4x-10

OpenStudy (anonymous):

Which are y=4x-1 y=4x-5 y=4x-10 y=4x+1

OpenStudy (anonymous):

That way is way easier thank you

OpenStudy (anonymous):

@chris00 yea that works too did not mean to interrupt you :)

OpenStudy (anonymous):

my way grasps the concept of a line visually. whereas a general formula is just a mere plug and chug. but no thats fine to interrupt. more ways the better

OpenStudy (anonymous):

graph it to check if (3,2) lies on the line though. so you can be sure that the line is correct